Tripling Money How long does it take for money to triple in an account paying 5% interest compounded annually?

Answer by stanbon(75887)

Tripling Money How long does it take for money to triple in an account paying 5% interest compounded annually?
----
Formula:: A(t) = P(1+(r/n))^(nt)
-----
Tripling means P becomes 3P
------
Solve:
3P = P(1+(0.05/1))^(1*t)
----
(1.05)^t = 3
----
t = log(3)/log(1.05)
----
t = 22.52 years
--------------------
Cheers,
Stan H.
-----------------